Using the Frobenius method, solve the following ODE:

x²y"+4xy'+(x²+2)y=0


1
Expert's answer
2021-05-24T17:30:51-0400

Let the solution of above differential equation be


"y=a_0x^m+a_1x^{m+1}+a_2x^{m+2}+...+a_nx^{m+n}"




"\dfrac{dy}{dx}=ma_0x^{m-1}+(m+1)a_1x^m"


"+(m+2)a_2x^{m+1}+...+(m+n)a_nx^{m+n-1}"


"\dfrac{d^2y}{dx^2}=m(m-1)a_0x^{m-2}+(m+1)ma_1x^{m-1}"


"+(m+2)(m+1)a_2x^{m}+..."

"+(m+n)(m+n-1)a_nx^{m+n-2}"


"x^2\dfrac{d^2y}{dx^2}+4x\dfrac{dy}{dx}+(x^2+2)y=0"


"x^2\big[m(m-1)a_0x^{m-2}+(m+1)ma_1x^{m-1}"

"+(m+2)(m+1)a_2x^{m}+..."

"+(m+n)(m+n-1)a_nx^{m+n-2}\big]"

"+4x\big[ma_0x^{m-1}+(m+1)a_1x^m"

"+(m+2)a_2x^{m+1}+...+(m+n)a_nx^{m+n-1}\big]"

"+(x^2+2)\big[a_0x^m+a_1x^{m+1}+a_2x^{m+2}+..."

"+a_nx^{m+n}\big]=0"

"(m^2+3m+2)a_0x^m+"

"+\displaystyle\sum_{n=1}^{\infin}((n+m)(n+m-1)+4(m+n)+2)a_nx^{m+n}"

"+\displaystyle\sum_{n=1}^{\infin}(n+m+1)a_{n-1}x^{m+n}=0"



Equating the coefficient of lowest degree in "x" (which is "x^m" ) equal to zero



"m^2+3m+2=0"


"m=-2,-1"

"m=-2"



"\displaystyle\sum_{n=1}^{\infin}((n-2)(n-3)+4(n-2)+2)a_n"

"+\displaystyle\sum_{n=1}^{\infin}(n-1)a_{n-1}x^{n-2}=0, n\geq1"

"a_n=-\dfrac{a_{n-1}}{n}, n\geq1"

"a_1=-a_0"

"a_2=-\dfrac{a_{1}}{2}=\dfrac{a_{0}}{2}"

"a_3=-\dfrac{a_{2}}{3}=-\dfrac{a_{0}}{6}"

"y_1(x)=a_0x^{-2}(1-x+\dfrac{x^2}{2}-\dfrac{x^3}{6}+...)"


"m=-1"



"\displaystyle\sum_{n=1}^{\infin}((n-1)(n-2)+4(n-1)+2)a_n"

"+\displaystyle\sum_{n=1}^{\infin}na_{n-1}=0, n\geq1"

"a_n=-\dfrac{a_{n-1}}{n+1}, n\geq1"

"a_1=-\dfrac{a_{0}}{2}"

"a_2=-\dfrac{a_{1}}{3}=\dfrac{a_{0}}{6}"

"a_3=-\dfrac{a_{2}}{4}=-\dfrac{a_{0}}{24}"

"y_2(x)=a_0x^{-1}(1-\dfrac{x}{2}+\dfrac{x^2}{6}-\dfrac{x^3}{24}+...)"




"y=C_3y_1(x)+C_4y_2(x)"


"=C_1x^{-2}(1-x+\dfrac{x^2}{2}-\dfrac{x^3}{6}+...)"

"+C_2x^{-1}(1-\dfrac{x}{2}+\dfrac{x^2}{6}-\dfrac{x^3}{24}+...)"